OpenStudy (karatechopper):

Describe how to write an algebraic expression from a real world situation. Include a definition of algebraic expression in your own words.

OpenStudy (chaise):

Three oranges and two apples are bought for $8. Eight oranges and four apples are bought for $21 Find the value of the oranges and the apples? Let o = oranges Let a = apples 3o + 2a = 8 8o + 4a = 21 You can solve these simultaneously. I'm not sure if it will work, but this is your example.

myininaya (myininaya):

Pretend we symbolize apple with A and coke cola with C and deodorant with D You go to the grocery store and you buy 2 apples and 3 coke colas so you have 2A+3C but you get to the car and realize your list said 5 apples and 6 coke colas and 1 deodorant so you need 5A+6C+1D so how do you figure whats left for you to get do what you need -what you got so we have (5A+6C+1D)-(2A+3C)=3A+3C+1D so this means you need 3 more apples and 3 more coke coloas and 1 deodorant

myininaya (myininaya):

chaise your problem sounds cool

OpenStudy (chaise):

Yours also. I think yours is easier for KarateChopper to understand though. Jumping straight into simultaneous equations probably wasn't the best idea.

OpenStudy (karatechopper):

thanks chaise and myininaya, good try though chaise's kinda helped but by the time it was up there i already figured mine out but it was good for reference. :)